OpenStudy (anonymous):

what is 3.25 with 5 repeating? will medal

OpenStudy (campbell_st):

what is the rest of the question...?

OpenStudy (anonymous):

into a fraction

OpenStudy (anonymous):

ss

OpenStudy (mathmate):

Multiply by 100 subtract 10 times the original number: 325.555555... -32.555555... =293 so 325.55555=293/(100-10)=293/90

OpenStudy (anonymous):

thank you

OpenStudy (mathmate):

You're welcome! :)